OpenStudy (anonymous):

Below are the steps to solve an equation: Step 1: |x - 4| + 1 = 8 Step 2: |x - 4| = 8 - 1 Step 3: |x - 4| = 7 Which of the following is a correct next step to solve the equation? x - 4 = -7 x + 4 = -7 -x - 4 = 7 x + 4 = 7

hartnn (hartnn):

if |a| = b then a= b or a = -b if |x-4| =7 then which 2 equation can you form ?

OpenStudy (anonymous):

7-4=x maybe?

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

let me give you an example :) if |y+9| = 10 then y+9 = 10 or y+9 = -10 so for |x-4| =7 what would it be ?

OpenStudy (anonymous):

x-4=-7 or x-4=7

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

that is correct! so which option ?

OpenStudy (anonymous):

x+4=7

OpenStudy (anonymous):

D

hartnn (hartnn):

nopes x-4 cannot change to x+4 see the first option x-4 = -7 :)

OpenStudy (anonymous):

ohhh :) A then thanks sooooo much

OpenStudy (anonymous):

can you help with one more ??

hartnn (hartnn):

welcome ^_^ ans sure :)

OpenStudy (anonymous):

The table shows the solution to the equation |2x - 3| - 1 = 2: Step 1 |2x - 3| = 2 + 1 Step 2 |2x - 3| = 3 Step 3 2x - 3 = 3 or 2x + 3 = 3 Step 4 2x = 6 or 2x = 0 Step 5 x = 3 or x = 0 Which is the first incorrect step? Step 1 Step 3 Step 5 Solution is correct

hartnn (hartnn):

after |2x - 3| = 3 we get 2x-3 =3 or 2x-3 = -3

hartnn (hartnn):

but step 3 shows something different

OpenStudy (anonymous):

soo step 3 is incorrect then right?

hartnn (hartnn):

yes!

OpenStudy (anonymous):

thats the first incorrect one then yayy :D thanks

OpenStudy (anonymous):

up for one more lol this is because I am trying to figure out all of these different concepts the teacher has introduced in multiple different assignments makes it confusing @hartnn

hartnn (hartnn):

ask in a different thread :) so that others can help u too!